Two balls are projected upward simultaneously with speed 40m/s and 60m/s . Relative position of second ball w.r.t first ball at time t=5s is

Grade:11

2 Answers

Hasan Naqvi
8 years ago
At t=5s;
Distance covered by first ball = 40*5 – ½ * 10 * 52 = 200 – 125 = 75m
Distance covered by second ball= 60 * 5 – ½ * 10 * 52 = 300 – 125 = 175m
 
Relative position of second ball wrt to first ball = 175-75 =100m
shubham
5 years ago
Range = velocity × time 
R1=40×5=200
R2=60×5=300
relative position is R2-R1 
so 300-200=100 
100 m is the  correct answer....